# Novel Pressure-Induced Transformations of PbTiO3

### v1

@@ Abstract
Perovskite oxides (ABO3) exhibit diverse functionalities ranging from superconductivity to ferroelectricity [1]. Lead titanate (PbTiO3), a classic ferroelectric, is widely studied for prototypical device applications [2, 3, 4, 5, 6, 7, 8, 9]. The symmetry breaking and loss of the inversion center, driven by Ti 3d3d–O $2p$ hybridization and Pb–O covalency [10, 11], result in a spontaneous polarization that stabilizes the tetragonal phase with a high Curie temperature of 763 K [12, 13, 14].
Compression of tetragonal (P4mm) PbTiO3 was first reported to suppress ferroelectricity near 12 GPa [15], consistent with a competition between long-range Coulomb interactions and short-range repulsion [11, 16, 17]. Subsequent studies combining diffraction and theoretical analysis revealed more complex behavior, including a pressure-induced morphotropic boundary at cryogenic temperatures that enhances the piezoelectric response [18, 19], as well as predictions of reentrant ferroelectricity [20]. Room-temperature Raman and X-ray studies confirmed the $P4mm\to I4/mcm$ transition near 18–20 GPa and suggested a recovery of a polar $I4cm$ state between 37–50 GPa based on improved fitting profiles [21].
@@
A recent theoretical and experimental investigation [22] extended these findings to higher pressures and identified a monoclinic $P2_{1}/m$ phase as the most stable structure of PbTiO3 above $\sim$80 GPa. This phase is centrosymmetric and nonpolar, consistent with second-harmonic generation (SHG) measurements performed as a function of pressure, which ruled out any polar response at both 300 K and 10 K. However, the stability and nature of the high-pressure structure of PbTiO3 under combined high-pressure and high-temperature conditions remain poorly constrained. Motivated by these results, we carried out high-pressure experiments on PbTiO3 up to 100 GPa and 1400 K using laser heating. Contrary to previous theoretical predictions that did not account for the possible dissociation of PbTiO3, the compound does not transform into the predicted monoclinic phase but instead decomposes into PbO and TiO2 at high temperature, and remains in the $I4/mcm$ phase under room-temperature compression. Through PXRD-assisted crystal structure predictions with XtalOpt [23, 24], we could identify, among the PbO products, the known $\alpha$ phase and a previously unreported $\delta$ phase. These findings reveal unexpected chemical reactivity and emphasize the kinetic and thermodynamic complexity of PbTiO3 under extreme PPTT conditions.
@@
It is useful to contrast the behavior of PbTiO3 with that of other oxide perovskites. Orthorhombic aluminate and stannate perovskites GdAlO3, YAlO3, and CaSnO3 exhibit the well known octahedral tilting on compression [55, 56, 57]. KNbO3, a prototypical ferroelectric perovskite, exhibits a sequence of pressure-induced phase transitions—from orthorhombic to tetragonal near 8 GPa, to cubic around 12 GPa, and finally to a distinct orthorhombic (GdFeO3-type) structure near 40 GPa [58, 59]. In contrast, CaTiO3, the mineral perovskite, remains stable on cold compression up to 60 GPa but transforms upon heating at higher pressure to a monoclinic CaTi2O5 phase and CaO [60, 61]. Of the magnetic perovskites, LaCrO3 undergoes a transition from rhombohedral to orthorhombic near 5 GPa [62, 63], whereas HoCrO3 transforms to a monoclinic phase above 20 GPa [64, 65]. Like PbTiO3, iron-rich (MgxFe1-x)SiO3 perovskite breaks down to form mixed oxides at lower mantle $P$-$T$ conditions similar to those explored here [66], whereas magnesium-rich (MgxFe1-x)SiO3 remains stable but transforms to a post-perovskite phase at higher pressures [67].
In conclusion, PbTiO3 exhibits a remarkable temperature-dependent response to compression. Two metastable pathways can be accessed on compression: one leads toward post-perovskite structures, whereas another preserves a nonpolar perovskite configuration that remains stable up to 100 GPa at room temperature  [22]. At elevated $P$-$T$ conditions, however, the system follows an equilibrium pathway that drives disproportionation into PbO and TiO2, producing a novel $\delta$-PbO polymorph distinct from the well-known $\alpha$-PbO (litharge) and $\beta$-PbO (massicot) structures. This new PbO polymorph is predicted to remain semiconducting to at least 100 GPa. These results illustrate how temperature can be used to control kinetic pathways for the synthesis of novel phases at high pressures: we demonstrate that in addition to the predicted reentrant ferroelectric and post-perovskite phases, PbTiO3 undergoes disproportionation, thereby expanding our understanding of the chemistry and diverse properties of oxide perovskites.
@@ Acknowledgments
This work was supported by the U.S. National Science Foundation (NSF) under Grants DMR-2119308 (HF, MA, RJH), DMR-2119065 (EZ), and PHY-2020249 (CMAP; SR). Additional support was provided by the U.S. Department of Energy (DOE), National Nuclear Security Administration (NNSA), through the Chicago/DOE Alliance Center (CDAC) under Award DE-NA0004153 (HF, MA, EZ, RJH), and by the DOE Office of Fusion Energy Sciences (OFES) under Award DE-SC0020340 (EZ, RH). We acknowledged the European Synchrotron Radiation Facility (ESRF) for provision of beamtime under proposal 5917. Portions of this work were performed at HPCAT (Sector 16), Advanced Photon Source (APS), Argonne National Laboratory (ANL), a DOE Office of Science (SC) user facility operated under Contract DE-AC02-06CH11357. HPCAT operations were supported by the DOE-NNSA Office of Experimental Sciences. Calculations were performed at the Center for Computational Research, University at Buffalo (http://hdl.handle.net/10477/79221). This work is a contribution of the George W. Crabtree Institute for Discovery and Sustainability.

### v2

# Novel Transformations of PbTiO3 with Pressure and Temperature

@@ Abstract
Perovskite oxides (ABO3) exhibit diverse functionalities ranging from superconductivity to ferroelectricity [85]. Lead titanate (PbTiO3), a classic ferroelectric, is widely studied for prototypical device applications [7, 70, 52, 82, 36, 86, 33, 54]. Early work by Shirane et al. established its ferroelectric phase transition near 760 K using X-ray and neutron diffraction [71], while Burns et al. demonstrated soft-mode behavior using Raman spectroscopy [13]. These lattice instabilities result in a spontaneous polarization that stabilizes the tetragonal $P4mm$ phase with a high Curie temperature of 763 K [71, 83, 57], above which the cubic $Pm\bar{3}m$ phase remains stable up to the melting temperature near 1558 K [23]. The seminal work by Cohen identified the role of Pb–O hybridization in stabilizing ferroelectricity [15, 17].
Subsequent high-pressure Raman studies by Sanjurjo et al. revealed strong pressure-induced phonon softening [67], consistent with a competition between long-range Coulomb interactions and short-range repulsion [17, 60, 14]. PbTiO3 later emerged as a key end member in relaxor ferroelectrics with giant electromechanical coupling [56]. At elevated pressures, increasingly complex behavior was reported, including pressure-induced morphotropic phase boundaries at cryogenic temperatures, where PbTiO3 transitions from tetragonal $P4mm$ to monoclinic and then rhombohedral symmetry, enhancing the piezoelectric response [81, 2]. These studies led to proposals of reentrant ferroelectricity, in which the nonpolar $I4/mcm$ phase transforms into the polar $I4cm$ structure above $\sim$30 GPa [38]. A later study showed that room-temperature Raman and X-ray measurements confirmed the $P4mm\rightarrow I4/mcm$ transition near 18–20 GPa and suggested a recovery of a polar $I4cm$ state between 37–50 GPa based on improved fitting profiles [32]. More recently, a combined theoretical and experimental investigation extended these studies to higher pressures and identified a centrosymmetric monoclinic $P2_{1}/m$ phase as the most stable structure of PbTiO3 above $\sim$80 GPa [16]. The centrosymmetric nature of this phase implies the absence of ferroelectricity and is consistent with second-harmonic generation measurements performed as a function of pressure, which ruled out any polar response at both 300 K and 10 K. This contrasting pressure response of ferroelectricity was highlighted in the classic work of Samara [66], who demonstrated that pressure can either suppress or enhance ferroelectric instabilities depending on whether they originate from zone-center or zone-boundary soft modes.
@@
However, the stability and nature of PbTiO3 under combined high $P$-$T$ conditions remain poorly constrained. Motivated by this unresolved question, we carried out high-pressure experiments on PbTiO3 up to 100 GPa and 1400 K using laser heating. Such high–PPTT diamond-anvil cell (DAC) experiments have previously enabled the discovery and synthesis of numerous novel functional materials, including transition-metal nitrides, polar ordered oxynitride perovskites, and superconducting superhydrides [84, 78, 72]. Until high pressure stability of perovskites was discovered, multicomponent oxides were thought to breakdown to simple oxides at high PPTT conditions. Contrary to theoretical predictions that did not account for the possible dissociation of PbTiO3, our experiments show that the compound does not transform into the predicted monoclinic phase at high temperature but instead decomposes into PbO and TiO2, while remaining in the $I4/mcm$ phase under room-temperature compression. Through PXRD-assisted crystal structure predictions with XtalOpt [63, 25], we identify both the known $\alpha$-PbO phase and a previously unreported $\delta$-PbO polymorph. Together, these results reveal unexpected chemical reactivity and underscore the kinetic and thermodynamic complexity of PbTiO3 under extreme pressure–temperature conditions.
@@
Direct experimental discrimination between metallic and semiconducting behavior in the high-pressure PbO polymorphs is not feasible due to the limited sample quantities, multiphase nature, and additional pressure-induced transformations of $\alpha$-PbO. Consequently, the electronic properties discussed above are assessed entirely through first-principles calculations.
@@
It is useful to contrast the behavior of PbTiO3 with that of other oxide perovskites. Orthorhombic aluminate and stannate perovskites GdAlO3, YAlO3, and CaSnO3 exhibit the well known octahedral tilting on compression [65, 64, 87]. KNbO3, a prototypical ferroelectric perovskite, exhibits a sequence of pressure-induced phase transitions—from orthorhombic to tetragonal near 8 GPa, to cubic around 12 GPa, and finally to a distinct orthorhombic (GdFeO3-type) structure near 40 GPa [62, 37]. In contrast, CaTiO3, the mineral perovskite, remains stable on cold compression up to 60 GPa but transforms upon heating at higher pressure to a monoclinic CaTi2O5 phase and CaO [77, 24]. Of the magnetic perovskites, LaCrO3 undergoes a transition from rhombohedral to orthorhombic near 5 GPa [88, 26], whereas HoCrO3 transforms to a monoclinic phase above 20 GPa [8, 45]. Although there is no evidence of PbTiO3 in the deep mantle, its behavior may be compared to that of the well studied (Mg,Fe)SiO3. Iron-rich (MgxFe1-x)SiO3 perovskite breaks down to form mixed oxides at lower mantle $P$-$T$ conditions similar to those explored here [27], whereas magnesium-rich (MgxFe1-x)SiO3 remains stable but transforms to a post-perovskite phase at higher pressures [30].
In conclusion, PbTiO3 exhibits a remarkable temperature-dependent response to compression. Two metastable pathways can be accessed on compression: one leads toward post-perovskite structures, whereas another preserves a nonpolar perovskite configuration that remains stable up to 100 GPa at room temperature  [16]. At elevated $P$-$T$ conditions, however, the system follows an equilibrium pathway that drives disproportionation into PbO and TiO2, producing a novel $\delta$-PbO polymorph distinct from the well-known $\alpha$-PbO (litharge) and $\beta$-PbO (massicot) structures. This new PbO polymorph is predicted to remain semiconducting to at least 100 GPa. These results illustrate how temperature can be used to control kinetic pathways for the synthesis of novel phases at high pressures: we demonstrate that in addition to the predicted reentrant ferroelectric and post-perovskite phases, PbTiO3 undergoes disproportionation, thereby expanding our understanding of the chemistry and diverse properties of oxide perovskites. Finally, we point out that early high-pressure studies suggested that all complex oxides ultimately disproportionate to form simple close packed oxide phases at sufficiently high pressures [9]. Later, static compression experiments demonstrated that a wide range of complex oxides instead stabilize dense crystal structures, notably perovskite [27] and later post-perovskite [30] phases. Our findings reveal a pathway for PbTiO3 that follows the original paradigm.
@@ Acknowledgments
This work was supported by the U.S. National Science Foundation (NSF) under Grants DMR-2119308 (HF, MA, RJH), DMR-2119065 (EZ), and PHY-2020249 (CMAP; SR). Additional support was provided by the U.S. Department of Energy (DOE), National Nuclear Security Administration (NNSA), through the Chicago/DOE Alliance Center (CDAC) under Award DE-NA0004153 (HF, MA, EZ, RJ), and by the DOE Office of Fusion Energy Sciences (OFES) under Award DE-SC0020340 (EZ, RH). We acknowledged the European Synchrotron Radiation Facility (ESRF) for provision of beamtime under proposal 5917. Portions of this work were performed at HPCAT (Sector 16), Advanced Photon Source (APS), Argonne National Laboratory (ANL), a DOE Office of Science (SC) user facility operated under Contract DE-AC02-06CH11357. HPCAT operations were supported by the DOE-NNSA Office of Experimental Sciences. Calculations were performed at the Center for Computational Research, University at Buffalo (http://hdl.handle.net/10477/79221). This work is a contribution of the George W. Crabtree Institute for Discovery and Sustainability.
@@
Figure S1: Simulated XRD patterns at 87 GPa over the laser-heated sample of PbTiO3.